show solution for (7x-3)+ 77+50= 180.
Answer by ewatrrr(24785)


Hi
solving (focusing on isolating x on one side of the equation)
1*(7x-3)+ 77+50= 180 |removing parenthesis
7x-3 + 77 + 50= 180 |combining Like Terms
7x + 124 = 180 |subtracting 124 from both sides of the equation
7x = 56 |Dividing both sides of the equation by 7 to finish!
x = 8